In a normal distribution with mu = 25 and sigma = 3, what number corresponds to z = -3?
Answer by jim_thompson5910(35256)

z = (x - mu)/(sigma)

-3 = (x - 25)/(3)

-3*3 = x - 25

-9 = x - 25

-9+25 = x

16 = x

x = 16

So the number x = 16 corresponds to z = -3